OpenStudy (anonymous):

Evaluate the integral: e^2x-1/e^x dx

OpenStudy (ipwnbunnies):

Separate the expression into two terms Pretty easy integral from there.

OpenStudy (anonymous):

Like 2^2x/e^x- 1/e^x ?

OpenStudy (zale101):

\[\int\limits_{}^{} \frac{e^{2x}-1}{e^x}~dx=\int\limits_{}^{}\frac{e^{2x}}{e^x}-\frac{1}{e^x}~dx=\int\limits_{}^{}\frac{e^{2x}}{e^x}~dx-\int\limits_{}^{}\frac{1}{e^x}~dx\]

OpenStudy (idku):

which is\[\int\limits_{ }^{ }e^x~dx~-\int\limits_{}^{ }~e^{-x}~dx\]

OpenStudy (zale101):

\[\int\limits_{}^{}\frac{e^{2x}}{e^x}~dx-\int\limits_{}^{}\frac{1}{e^x}~dx\] For this part \[\int\limits_{}^{}\frac{e^{2x}}{e^x}~dx\] it can also be writing as \[\int\limits_{}^{}\frac{e^{x^2}}{e^x}~dx\] the e^x cancels out, and you'll get \[\int\limits_{}^{}e^x~dx\]

OpenStudy (zale101):

for the second part, idku use this method \(\frac{1}{x}=x^{-1}\) from algebra \[\int\limits_{}^{}e^x~dx-\int\limits_{}^{}e^{-x}~dx\]

OpenStudy (zale101):

my latex was wrong a lil bit \(\Large e^{2x}=e^{x+x}=e^{x}*e^{x}=(e^{x})^2\)